OpenStudy (anonymous):

Find parametric equation for rectangular equation: (y+3)^2 = 4(x-2)

OpenStudy (anonymous):

(y+3)^2+8 ----------- =x 4 (y+3)=t y= t-3 x= t^2+8 -------- 4

OpenStudy (anonymous):

should be sideway parabola

OpenStudy (anonymous):

you will see that they are same http://www.wolframalpha.com/input/?i=Plot%5B%28y%2B3%29%5E2+%3D+4%28x-2%29%2C%7Bx%2C0%2C6%7D%2C%7By%2C-6%2C1%7D%5D ParamtricPlot{(t^2+8)/4,t-3}

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I don't get this part x= t^2+8 -------- 4

OpenStudy (anonymous):

I set t to be t=y+3 and plug into here (y+3)^2+8 ----------- =x 4

OpenStudy (anonymous):

ok, i got it